I need help solving these constraints
3A-2B>=0
2A-1B<=200

Answer by ankor@dixie-net.com(22740)

I need help solving these constraints
3A-2B>=0
2A-1B<=200
:
Multiply the 2nd equation by -2 and add the 1st equation to it:
(This also reverses the inequality sign making them the same
-4a + 2b >= -400
3a - 2b >= 0
-------------- adding eliminates b
-1a + 0b >= -400
:
Multiply equation by -1 to get a positive a, also reverses the inequality sign
a <= +400
:
Substitute 400 for a in the 1st equation:
3(400) - 2b >= 0
1200 - 2b >= 0
-2b >= -1200
Multiply equation by -1 to get a positive b, reverses the inequality sign again
b <= 1200
b <= 1200/2
b <= 600
:
you can plot a as vertical line at x=400 and b as a horizontal line at y=600
The area of feasibility; at or below y = 600 and, at or to the left of x = 400
:
Did this help?